OpenStudy (anonymous):

graph y = x^3 – 4x^2 + 6. Describe what you see. How is what you observe here different from what you observed in part A?

OpenStudy (anonymous):

PART A is :A.Graph y = x^3 – 2x^2 + 6. Describe what you see.

OpenStudy (anonymous):

yal still there ?

OpenStudy (anonymous):

in first equation you can set x = 0 to get y = 6

OpenStudy (anonymous):

but you can't sub 6 to get the same answer ...

OpenStudy (anonymous):

the answer would be : 6 = = x^3 – 4x^2 + 6 0=x^3-4x^2 either x = zero or .. x=4

OpenStudy (anonymous):

no the same matter of equation 2 in part A